# Simulation-based inference of galaxy properties from JWST pixels

### v1

@@
Patricia Iglesias-Navarro 1122    Marc Huertas-Company 11223344    Pablo Pérez-González 55    Johan H. Knapen 1122    ChangHoon Hahn 77    Anton M. Koekemoer 66    Steven L. Finkelstein 88    Natalia Villanueva 88    Andrés Asensio Ramos 1122
The spectral energy distributions (SEDs) of galaxies offer detailed insights into their stellar populations, capturing key physical properties such as stellar mass, star formation history (SFH), metallicity, and dust attenuation. However, inferring these properties from SEDs is a highly degenerate inverse problem, particularly when using integrated observations and a limited number of photometric bands. We present an efficient Bayesian SED-fitting framework tailored to multiwavelength pixel photometry from the JWST Advanced Deep Extragalactic Survey (JADES). Our method employs simulation-based inference to enable rapid posterior sampling across galaxy pixels, leveraging the unprecedented spatial resolution, wavelength coverage, and depth provided by the survey. It is trained on synthetic photometry generated from MILES stellar population models, incorporating both parametric and non-parametric SFHs, realistic noise, and JADES-like filter sensitivity thresholds. We validate this amortised inference approach on mock datasets, achieving robust and well-calibrated posterior distributions, with an $R^{2}$ score of 0.99 for stellar mass. Applying our pipeline to real observations, we derive spatially resolved maps of stellar population properties down to $\mathrm{S/N}_{\rm{pixel}}=5$ (averaged over F277W, F356W, F [^4] W) for $1083$ JADES galaxies and $\sim$2 million pixels with spectroscopic redshifts. These maps enable the identification of dusty or starburst regions and offer insights into mass growth and the structural assembly. We assess the outshining phenomenon by comparing pixel-based and integrated stellar mass estimates, finding limited impact only in low-mass galaxies ($<10^{8}\,M_{\odot}$) but systematic differences of $\sim$0.20 dex linked to SFH priors. With an average posterior sampling speed of $10^{-4}$ seconds per pixel and a total inference time of $\sim$1 CPU-day for the full dataset, our model offers a scalable solution for extracting high-fidelity stellar population properties from HST+JWST datasets, opening the way for statistical studies at sub-galactic scales.
@@ 1 Introduction
The Hubble Space Telescope (HST) and the James Webb Space Telescope (JWST) have provided high-resolution, multi-wavelength spectral energy distributions (SEDs) of hundreds of thousands of galaxies. This allows us for the first time to analyse their stellar population properties per pixel, not only for local sources, but even for moderately high redshifts $(z\sim 2-10)$ (Pérez-González et al., 2023; Watson et al., 2024; Lines et al., 2025; Gibson et al., 2024; Giménez-Arteaga et al., 2024; Harvey et al., 2025), opening a window to study the processes of galactic structure, growth and evolution in significant detail.
@@
This vast amount of high-resolution data continues to grow, while the methods we use to analyse it have largely remained the same. Traditional SED analysis techniques are computationally intensive and often prohibitively slow. Standard optimization approaches (e.g., Sawicki & Yee, 1998; Heavens et al., 2000; Tojeiro et al., 2007; Dye, 2008; Iyer & Gawiser, 2017) may take minutes per galaxy, while Bayesian methods (e.g., Pacifici et al., 2012; Carnall et al., 2018; Boquien et al., 2019; Johnson et al., 2021a), critical for accounting for uncertainties and degeneracies in stellar properties, can demand up to 10-100 CPU hours per galaxy (Tacchella et al., 2022). This bottleneck in computational efficiency prevents a scalable analysis of galaxy stellar populations.
@@
A solution to this challenge lies in combining Bayesian inference with machine learning tools to achieve rapid and accurate posterior distributions for galaxy physical properties. In particular, simulation-based inference (SBI; Cranmer et al., 2020) has emerged as a powerful framework for learning complex, non-parametric probability distributions from synthetic data (e.g., Hahn & Melchior, 2022; Khullar et al., 2022; Li et al., 2023; Iglesias-Navarro et al., 2024; Kwon & Hahn, 2024). This method enables an amortised inference: after an initial training phase on synthetic galaxy samples, the model can estimate posterior distributions for new observations in a fraction of the traditional time, bypassing the need for retraining. This approach captures degeneracies among parameters and their uncertainties effectively, even in the context of the sparse photometric data typical of deep-field surveys.
@@
In this paper we extend this novel and flexible approach to estimate the stellar population properties of resolved galaxies in the Hubble Ultra Deep Field (HUDF; Beckwith et al., 2006). Using filters in the Advanced Camera for Surveys on the HST (ACS; Ryon, 2023) and the Near Infrared Camera on the JWST (NIRCam; Rieke et al., 2023a), our method reconstructs distributions of stellar mass, star formation rate, metallicity, and dust content per pixel for thousands of galaxies. Special attention is given to the uncertainties in flux measurements, depth variations, and PSF matching between the HST and JWST filters, optimising the training to reflect realistic observational conditions. Our model is applicable to galaxies across a range of morphologies and redshifts up to $z\sim 10$, offering a comprehensive framework for studying stellar population properties across cosmic time.
@@
We obtain high-resolution, two-dimensional maps of stellar population properties of over $1000$ galaxies in the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al., 2023b). The pixel-by-pixel precision achieved by this approach enables us to identify and analyse regions of dusty starbursts, uncover clues about their mass growth and structural assembly processes, and gain broader insights into the star formation in galaxies during the earliest cosmic epochs.
@@
The outline of the paper is as follows: In Sect. 2 we detail the observational data. In Sect. 3 we describe the Bayesian inference framework, using SPS (Sect. 3.1), and a neural density estimator (Sect. 3.2). In Sect. 4 we test the model with mock observations, and in Sect.  5 with individual pixels from the JADES survey. We then perform a pixel-by-pixel analysis of the full dataset in Sect. 6. Finally, in Sect. 7, we discuss our model’s inferred galaxy properties and the future applications of this technique.
We assume a flat cosmology with $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$, and a Hubble constant $H_{0}=70$ km s-1Mpc-1. We use AB magnitudes (Oke & Gunn, 1983) and a Chabrier initial mass function (Chabrier, 2003).
@@ 2 Observational data
We selected a sample of $1083$ galaxies in the HUDF (Beckwith et al., 2006) within the GOODS-South field (GOODS-S; Giavalisco et al., 2004), where the JADES (Eisenstein et al., 2023b), JEMS (Williams et al., 2023) and FRESCO (Oesch et al., 2023) surveys overlap. The MUSE (Bacon et al., 2017, 2023) survey also covers this region, which guarantees spectroscopic redshifts for a large fraction of the galaxies. We used no particular criterion but selected all galaxies from the JADES segmentation map (Rieke et al., 2023b; D’Eugenio et al., 2024) in a region where all the $19$ filters included in Table 1, as well as spectroscopic redshifts, are available. Later in the inference step we did use a SNR criterion, as described in Sect. 5. By relying only on spectroscopic redshifts we reduce a major uncertainty of the SED fitting, namely the uncertainty of the photometric redshift.
@@
We use photometry from the HST ACS filters and JWST NIRCam wide-band and medium-band filters (see Table 1). This provides a wide wavelength coverage from 0.4 $\mu$m to 4.8 $\mu$m, from rest-frame ultra-violet and optical to near-infrared light. The HST mosaics include data from GOODS (Giavalisco et al., 2004) that overlap with imaging data from the HUDF (Beckwith et al., 2006; Ellis et al., 2013; Illingworth et al., 2013; Koekemoer et al., 2013) and CANDELS (Grogin et al., 2011; Koekemoer et al., 2011), combined with other archival HST imaging data, which have all been reprocessed using up-to-date calibrations to produce mosaics at a scale of ${\prime\prime}$03 per pixel, astrometrically aligned to Gaia-DR3111https://www.cosmos.esa.int/web/gaia/dr3. For the JADES mosaics we use the DR2 of JADES of the GOODS field, which are described in more detail by the JADES team (Rieke et al., 2023b; Eisenstein et al., 2023a, b).
We subtracted the background signal using the Background2D function from Photutils (Bradley et al., 2024). We estimated the background by first dividing the mosaics into grids of $200\times 200$ pixels $(0.1^{\prime}\times 0.1^{\prime})$. We then compute, for each grid, the background level using the $\sigma$-clipping method, averaging the flux of pixels whose signal was not above $3\sigma$ of the total flux distribution in each grid.
We convolved all mosaics with the point spread function (PSF) of the reddest wide-band filter: F [^4] W, except for the F460M and F480M mosaics, that were left unconvolved. We built the PSFs using Photutils from a list of unresolved objects in the HUDF (Pirzkal et al., 2005). We generated the kernels with the function Create_matching_kernel, again from the Photutils package, using a CosineBellWindow filter with $\alpha=0.5$. We normalise the kernels so they did not affect the average flux values.
@@ 3 Simulation-based inference
We aim to infer the posterior distribution of galaxy properties, denoted as $\theta$, based on photometric observations $\{X_{i}\}$. Traditional Bayesian inference methods depend on the likelihood function $P(X|\theta)$, which describes the probability of observing the data given the parameters, to compute the posterior distributions through the Bayes’ theorem. However, such calculations become challenging in high-dimensional systems with complex degeneracies between those parameters. To address this issue, we employ SBI, which utilises forward models to simulate data from physical parameters and produce pairs of parameters and data points $(\theta,X^{\prime})$ through a forward model. By comparing the synthetic data points, $X^{\prime}$, with the actual observations, $X$, SBI enables the estimation of the posterior distribution of the parameters.
@@ 3.1 Forward model
We first generated a dataset of simulated composite stellar populations (CSP; interpretable as galaxies or galaxy pixels) with known ACS and NIRCam photometry, associated errors, and properties $\theta$. We relied on the stellar population synthesis (SPS) approach: starting from a library of empirical stellar spectra and assuming an initial mass function (IMF) and a set of isochrones, we retrieved the SEDs of simple stellar populations (SSPs) across varying ages and metallicities. We combined these using star formation histories (SFHs), considering constant metallicity, and applied a dust attenuation law. For a more detailed review, we refer the reader to Conroy (2013) and Iyer et al. (2025).
We used the pipeline provided by FSPS (Conroy et al., 2009; Conroy & Gunn, 2010) and Dense Basis (Iyer & Gawiser, 2017; Iyer et al., 2019), with MIST isochrones (Choi et al., 2016), the MILES stellar spectral library (Vazdekis et al., 2010), and the DL07 dust emission library (Draine & Li, 2007), as well as a Chabrier IMF (Chabrier, 2003). We considered a Calzetti dust attenuation law (Calzetti et al., 2000), coupling the birth-cloud attenuation to that from older stars, and generated both nebular continuum and emission using CLOUDY (Ferland et al., 2013, 2017), with gas-phase metallicity set equal to stellar metallicity and held constant throughout the galaxy’s lifetime. We used two sets of simulations with different SFHs. First, we considered a parametric tau-delayed model:
@@
where $t_{i}$ is the cosmic time at which the star formation begins, $\tau$ is the characterisic timescale for the exponential decline, and $t$ is the current cosmic time. For $t<t_{i},\;\operatorname{SFR}(t)=0$, as no star formation occurs before $t_{i}$.
We also constructed non-parametric SFHs with the Dense Basis module GP-SFH (Iyer & Gawiser, 2017; Iyer et al., 2019), allowing for complex behaviours like rejuvenation events, bursts, or sudden quenching without relying on a fixed functional form. Our selected prior assumes that the specific star formation rate (sSFR), i.e., the star formation rate normalized by stellar mass, across three equally spaced time bins, follows a Dirichlet distribution with a concentration parameter $\alpha=1$. This prior for SFHs was studied in detail by Leja et al. (2019). We also included the AGB circumstellar dust model from Villaume et al. (2015) and accounted for mass loss in our total calculations.
The priors for the parameters that control these components for both simulations are specified in Table 2. Rather than matching the observational distributions, we opted to keep the priors as uniform as possible, as inference with unbalanced priors is more prone to systematic effects and biases (Hahn et al., 2023). The exception is the redshift distribution, where although uniform, the ranges were chosen according to the dataset described in Sect. 2. We set a lower limit of $\log_{10}(M_{*}/\rm{M}_{\odot})=4.0$ for the prior for the stellar mass. At these very low stellar masses per pixel, SSP models become unreliable due to statistical fluctuations in the stellar population, particularly from stochastic sampling of the initial mass function (IMF, Cerviño & Valls-Gabaud 2003). This limitation is especially relevant in the outskirts of low-redshift galaxies, where the surface mass density can be extremely low. In these regions, inferred stellar population properties such as age and metallicity are highly uncertain. Therefore, instead of fitting each pixel independently, more robust strategies like binning adjacent pixels or analysing radial profiles should be used to improve the signal and reduce stochastic noise.
Table 2: Priors for the parameters of the simulation using the $\tau$-delayed and Dirichlet models.
@@
$\log_{10}(M_{*}^{\rm{formed}}/\rm{M}_{\odot})^{\dagger}$ U$(4.0,12.0)$ U$(4.0,12.0)$
$\tau$ [Gyr] LU$(10^{-2},100)$
$t_{i}$ [Gyr] U$(0,\mathrm{age}_{\mathrm{Universe}}(z))$
$\log_{10}$(sSFR) U$(-11.5,-7.5)$
$t_{25\%,50\%,75\%}$†† Dir$(\alpha=1.0,N=3)$
$[M/\mathrm{H}]$ U$(-2.3,0.4)$ U$(-2.3,0.4)$
$A_{V}$ U$(0.0,4.0)$ U$(0.0,4.0)$
$z$ U$(0.0,7.5)$ U$(0.0,7.5)$
@@
log($U$) $-2.0$
 [^2] For the Dirichlet SFHs, we selected the prior on $\log_{10}$(sSFR/(yr-1)) but inferred $\log_{10}$(SFR/($\rm{M}_{\odot}$yr-1)) instead. Fixed parameters include dust emission, nebular continuum, and emission information. U: Uniform, LU: Log-uniform, Dir: Dirichlet.
@@
  • $\dagger$ Formed stellar mass (integral of SFH). We also inferred surviving stellar mass.
[…]
@@
The output of the forward model were the fluxes in the 19 filters included in Table 1, including ACS filters and NIRCam wide-band and medium-band filters. To simulate realistic observational conditions, we modelled the flux uncertainties as a function of flux in each filter. We used the JADES segmentation map to identify all pixels classified as galaxies, and for each filter, we binned the pixel fluxes into five flux bins. Within each bin, we collected the flux uncertainties of the real observations and fit their distribution with a Gaussian. These fitted Gaussians, capturing the empirical uncertainty–flux relation, were then used to add noise to the simulated fluxes: for each simulated SED, the noise was sampled from the appropriate Gaussian based on its flux in each filter. This approach captures a broad range of uncertainties consistent with the observational data. Following the procedure described in Hahn & Melchior (2022), the network inputs included the fluxes with noise sampled from the distributions, as well as the standard deviations of the Gaussian distributions, to condition the inference by producing noise-aware posteriors.
@@
We transformed both fluxes and standard deviations to AB magnitudes, avoiding numerical instabilities in the training process caused by the wide dynamic range of possible flux values. We also took into account the $1\sigma$ depth limit in each filter $m_{i}^{b}$. We use the values obtained in Hainline et al. (2024) for the JADES mosaics. Although we repeated the calculations of the $1\sigma$ depth limits in the region were the galaxies were selected, we found negligible differences. To make this process amortised we use a mean value for the depth limit, not dependant of the specific regions where the galaxies are located. If the simulated magnitude in a filter $m_{i}$ for a CSP is fainter than that depth limit ($m_{i}>m_{\sigma_{i}}^{b}$), we transformed it to $m_{i}^{\prime}=100$ and and the noise in the input as $m_{\sigma_{i}}^{b}$. While $m_{i}^{\prime}=100$ is not physically meaningful (since it should be infinite), it is a convenient choice that avoids convergence problems in the calculations and allows clear identification of dropouts, cases where the signal is not detected in a given filter. These dropouts typically occur at bluer wavelengths than the Lyman break for high-z galaxies, although non-detections can also occur in redder filters, especially in low SNR photometry. Identifying these dropouts could also ensure a correct redshift estimate, although we have not addressed it in this work.
We obtained posterior distributions from the photometry to constrain the stellar population parameters included in Table 2 for CSPs, using the neural density estimator as explained below.

### v3

@@
Patricia Iglesias-Navarro    Marc Huertas-Company    Pablo Pérez-González    Johan H. Knapen    ChangHoon Hahn    Anton M. Koekemoer    Steven L. Finkelstein    Natalia Villanueva    Andrés Asensio Ramos
The spectral energy distributions (SEDs) of galaxies offer detailed insights into their stellar populations, capturing key physical properties such as stellar mass, star formation history (SFH), metallicity, and dust attenuation. However, inferring these properties from SEDs is a highly degenerate inverse problem, particularly when using integrated observations across a limited range of photometric bands. We present an efficient Bayesian SED-fitting framework tailored to multiwavelength pixel photometry from the JWST Advanced Deep Extragalactic Survey (JADES). Our method employs simulation-based inference to enable rapid posterior sampling across galaxy pixels, leveraging the unprecedented spatial resolution, wavelength coverage, and depth provided by the survey. It is trained on synthetic photometry generated from MILES stellar population models, incorporating both parametric and non-parametric SFHs, realistic noise, and JADES-like filter sensitivity thresholds. We validated this amortised inference approach on mock datasets, achieving robust and well-calibrated posterior distributions, with an $R^{2}$ score of 0.99 for stellar mass. Applying our pipeline to real observations, we derived spatially resolved maps of stellar population properties down to $\mathrm{S/N}_{\rm{pixel}}=5$ (averaged over F277W, F356W, and F [^4] W) for $1083$ JADES galaxies and $\sim$2 million pixels with spectroscopic redshifts. These maps enable the identification of dusty or starburst regions, offering insights into mass growth and structural assembly. We assessed the outshining phenomenon by comparing pixel-based and integrated stellar mass estimates, finding a limited impact only in low-mass galaxies ($10^{8}\,M_{\odot}$), but with systematic differences of $\sim$0.20 dex linked to SFH priors. With an average posterior sampling speed of $10^{-4}$ seconds per pixel and a total inference time of $\sim$1 CPU-day for the full dataset, our model offers a scalable solution for extracting high-fidelity stellar population properties from HST+JWST datasets, paving the way for statistical studies on sub-galactic scales.
@@ 1 Introduction
Spatially resolved spectral energy distribution (SED) fitting boasts a long history of use as a powerful tool for understanding internal galaxy structure and evolution. Early works by Abraham et al. (1999) and Zibetti et al. (2009) demonstrated how spatial variations in stellar populations and mass-to-light ratios can impact global galaxy properties. With high-resolution multi-wavelength imaging from the Hubble Space Telescope (HST), pixel-by-pixel SED fitting techniques were further developed (e.g. Wuyts et al. (2012), Lanyon-Foster et al. (2012)) revealing radial gradients in star formation and dust attenuation at intermediate redshifts ($z\sim 1$–2). These studies laid the groundwork for understanding internal galaxy diversity beyond integrated light measurements. A key challenge in performing unresolved SED fittings is the outshining effect, where bright, young star-forming regions dominate the observed light and mask older, more massive, but less luminous stellar populations, leading to systematic underestimations of total stellar mass (Sorba & Sawicki, 2015, 2018; Abdurro’uf & Akiyama, 2017; Smith & Hayward, 2018). This effect is particularly significant in galaxies with high specific star formation rates (sSFRs) and bursty star formation histories (SFHs).
Recent advances with HST and especially the James Webb Space Telescope (JWST) have transformed this field by enabling high-resolution, multi-band SEDs for hundreds of thousands of galaxies, extending pixel-level analyses to high redshifts ($z\sim 2$–10) (Pérez-González et al., 2023; Watson et al., 2025; Lines et al., 2025; Gibson et al., 2024; Giménez-Arteaga et al., 2024; Harvey et al., 2025). These data have revealed that outshining can cause up to 1 dex discrepancies between integrated and resolved stellar masses in low-mass, high-sSFR galaxies (Giménez-Arteaga et al., 2023, 2024; Fujimoto et al., 2025), while more massive galaxies show smaller offsets likely due to a smoother SFHs (Lines et al., 2025; Pérez-González et al., 2023). Theoretical studies support these observations, with hydrodynamic simulations demonstrating how bursty SFHs cause mass estimate biases (Narayanan et al., 2024), showing that resolved fitting recovers stellar masses well when deep multiband photometry is available (Cochrane et al., 2025; Mosleh et al., 2025). Model assumptions, priors, and spatial resolution further influence the interpretation of resolved SED fits (Harvey et al., 2025). Together, these observational and theoretical developments, powered by HST and JWST data, open up a new window onto galaxy evolution by enabling detailed, spatially resolved stellar population analyses over a wide range of cosmic timescales.
@@
This vast amount of high-resolution data continues to grow, while the methods we use to analyse it have largely remained the same. Traditional SED analysis techniques are computationally intensive and often prohibitively slow. Standard optimisation approaches (e.g. Sawicki & Yee, 1998; Heavens et al., 2000; Tojeiro et al., 2007; Dye, 2008; Iyer & Gawiser, 2017) can take several minutes per galaxy, while Bayesian methods (e.g. Pacifici et al., 2012; Carnall et al., 2018; Boquien et al., 2019; Johnson et al., 2021a) that are critical for accounting for uncertainties and degeneracies in stellar properties can demand up to 10-100 CPU hours per galaxy (Tacchella et al., 2022). This bottleneck in computational efficiency stands in the way of a scalable analysis of galaxy stellar populations.
@@
A solution to this challenge lies in combining Bayesian inference with machine learning tools to achieve rapid and accurate posterior distributions for galaxy physical properties. In particular, simulation-based inference (SBI; Cranmer et al., 2020) has emerged as a powerful framework for learning complex, non-parametric probability distributions from synthetic data (e.g. Hahn & Melchior, 2022; Khullar et al., 2022; Li et al., 2023; Iglesias-Navarro et al., 2024; Kwon & Hahn, 2024). This method enables an amortised inference: after an initial training phase on synthetic galaxy samples, the model can estimate posterior distributions for new observations in a fraction of the traditional time, bypassing the need for retraining. This approach captures degeneracies among parameters and their uncertainties effectively, even in the context of the sparse photometric data typical of deep-field surveys.
@@
In this paper, we extend this novel and flexible approach to estimate the stellar population properties of resolved galaxies in the Hubble Ultra Deep Field (HUDF; Beckwith et al., 2006). Using filters in the Advanced Camera for Surveys on the HST (ACS; Ryon, 2023) and the Near Infrared Camera on the JWST (NIRCam; Rieke et al., 2023a), our method reconstructs distributions of stellar mass, SFR, metallicity, and dust content per pixel for thousands of galaxies. Special attention is given to the uncertainties in flux measurements, depth variations, and PSF matching between the HST and JWST filters, optimising the training to reflect realistic observational conditions. Our model is applicable to galaxies across a range of morphologies and redshifts up to $z\sim 10$, offering a comprehensive framework for studying stellar population properties across cosmic time.
@@
We obtain high-resolution, two-dimensional maps of stellar population properties of over $1000$ galaxies in the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al., 2023b). The pixel-by-pixel precision achieved by this approach enables us to identify and analyse regions of dusty starbursts, uncovering clues about their mass growth and structural assembly processes and gaining broader insights into the star formation in galaxies during the earliest cosmic epochs.
@@
The outline of the paper is as follows. In Sect. 2, we detail the observational data. In Sect. 3, we describe the Bayesian inference framework, using stellar population synthesis (Sect. 3.1) and a neural density estimator (Sect. 3.2). In Sect. 4, we test the model with mock observations and then with individual pixels from the JADES survey Sect. 5. We describe our pixel-by-pixel analysis of the full dataset in Sect. 6. Finally, in Sects. 7 and 8, we discuss our model’s inferred galaxy properties and potential future applications of this technique. We assumed a flat cosmology with $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$, and a Hubble constant $H_{0}=70$ km s-1Mpc-1. We used AB magnitudes (Oke & Gunn, 1983) and a Chabrier initial mass function (Chabrier, 2003).
@@ 2 Observational data
We selected a sample of $1083$ galaxies in the HUDF (Beckwith et al., 2006) within the GOODS-South field (GOODS-S; Giavalisco et al., 2004), where the JADES (Eisenstein et al., 2023b), JEMS (Williams et al., 2023), and FRESCO (Oesch et al., 2023) surveys overlap. The MUSE (Bacon et al., 2017, 2023) survey also covers this region, which guarantees spectroscopic redshifts for a large fraction of the galaxies. We used no particular criterion and opted to select all the galaxies from the JADES segmentation map (Rieke et al., 2023b; D’Eugenio et al., 2024) in a region with all the available $19$ filters included in Table 1, as well as the spectroscopic redshifts. Later, in the inference step, we used a signal-to-noise ratio (S/N) criterion, as described in Sect. 5. By relying only on spectroscopic redshifts we reduce one major uncertainty of the SED fitting, namely, the uncertainty on the photometric redshift.
@@
[…]
We subtracted the background signal using the Background2D function from Photutils (Bradley et al., 2024). We estimated the background by first dividing the mosaics into grids of $200\times 200$ pixels $(0.1^{\prime}\times 0.1^{\prime})$. We then computed, for each grid, the background level using the $\sigma$-clipping method, averaging the flux of pixels whose signal was not above $3\sigma$ of the total flux distribution in each grid.
We convolved all the mosaics with the point spread function (PSF) of the reddest wide-band filter: F [^4] W, except for the F460M and F480M mosaics, which were left unconvolved. We built the PSFs using Photutils from a list of unresolved objects in the HUDF (Pirzkal et al., 2005). We generated the kernels with the function Create_matching_kernel, again from the Photutils package, using a CosineBellWindow filter with $\alpha=0.5$. We normalised the kernels so that they would not affect the average flux values.
@@ 3 Simulation-based inference
Our approach is aimed at inferring the posterior distribution of galaxy properties, denoted as $\theta$, based on photometric observations, $\{X_{i}\}$. Traditional Bayesian inference methods depend on the likelihood function $P(X|\theta)$, which describes the probability of observing the data given the parameters, to compute the posterior distributions through the Bayes’ theorem. However, such calculations become challenging in high-dimensional systems with complex degeneracies between those parameters. To address this issue, we employ SBI, which utilises forward models to simulate data from physical parameters and produce pairs of parameters and data points $(\theta,X^{\prime})$ via a forward model. By comparing the synthetic data points, $X^{\prime}$, with the actual observations, $X$, SBI enables the estimation of the posterior distribution of the parameters.
@@ 3.1 Forward model
First, we generated a dataset of simulated composite stellar populations (CSP; interpretable as galaxies or galaxy pixels) with known ACS and NIRCam photometry, associated errors, and properties, $\theta$. We relied on the stellar population synthesis (SPS) approach: starting from a library of empirical stellar spectra and assuming an initial mass function (IMF) and a set of isochrones, we retrieved the SEDs of simple stellar populations (SSPs) across varying ages and metallicities. We combined these using SFHs, assuming a constant metallicity, and we applied a dust attenuation law. For a more detailed review, we refer to Conroy (2013) and Iyer et al. (2025).
We used the pipeline provided by FSPS (Conroy et al., 2009; Conroy & Gunn, 2010) and Dense Basis (Iyer & Gawiser, 2017; Iyer et al., 2019), with MIST isochrones (Choi et al., 2016), the MILES stellar spectral library (Vazdekis et al., 2010), and the DL07 dust emission library (Draine & Li, 2007), as well as a Chabrier IMF (Chabrier, 2003). We considered a Calzetti dust attenuation law (Calzetti et al., 2000), coupling the birth-cloud attenuation to that from older stars, and we generated both nebular continuum and emission using CLOUDY (Ferland et al., 2013, 2017), with the gas-phase metallicity set equal to stellar metallicity and held constant throughout the galaxy’s lifetime. We used two sets of simulations with different SFHs. First, we considered a parametric tau-delayed model,
@@
where $t_{i}$ is the cosmic time at which the star formation begins, $\tau$ is the characterisic timescale for the exponential decline, and $t$ is the current cosmic time. For $t<t_{i},$ we have $\operatorname{SFR}(t)=0$, since no star formation occurs before $t_{i}$.
We also constructed non-parametric SFHs with the Dense Basis module GP-SFH (Iyer & Gawiser, 2017; Iyer et al., 2019), allowing for complex behaviours such as rejuvenation events, bursts, or sudden quenching without relying on a fixed functional form. Our selected prior assumes that the sSFR (i.e. the SFR normalised by stellar mass) across three equally spaced time bins, follows a Dirichlet distribution with a concentration parameter of $\alpha=1$. This SFH prior was studied in detail by Leja et al. (2019). Here, we also included the AGB circumstellar dust model from Villaume et al. (2015) and accounted for mass loss in our total calculations.
The priors for the parameters that control these components for both simulations are specified in Table 2. Rather than matching the observational distributions, we opted to keep the priors as uniform as possible, as inference with unbalanced priors is more prone to systematic effects and biases (Hahn et al., 2023). The exception is the redshift distribution, which remains uniform but whose ranges were chosen according to the dataset described in Sect. 2. We set a lower limit of $\log_{10}(M_{*}/\rm{M}_{\odot})=4.0$ for the prior for the stellar mass. At these very low stellar masses per pixel, SSP models become unreliable due to statistical fluctuations in the stellar population, particularly from stochastic sampling of the initial mass function (IMF, Cerviño & Valls-Gabaud 2003). This limitation is especially relevant in the outskirts of low-redshift galaxies, where the surface mass density can be extremely low. In these regions, inferred stellar population properties such as age and metallicity, are highly uncertain. Therefore, instead of fitting each pixel independently, more robust strategies such as binning adjacent pixels or analysing radial profiles should be used to improve the signal and reduce stochastic noise.
Table 2: Priors for the parameters of the simulation using the $\tau$-delayed and Dirichlet models. [^2] For the Dirichlet SFHs, we selected the prior on $\log_{10}$(sSFR/(yr-1)), but we inferred $\log_{10}$(SFR/($\rm{M}_{\odot}$yr-1)) instead. Fixed parameters include dust emission, nebular continuum, and emission information. U: Uniform, LU: Log-uniform, and Dir: Dirichlet.
@@
$\log_{10}(M_{*}^{\rm{formed}}/\rm{M}_{\odot})^{\dagger}$ U(4.0,12.0)(4.0,12.0) U(4.0,12.0)(4.0,12.0)
[…]
$\log_{10}$(sSFR) U(11.5,7.5)(-11.5,-7.5)
[…]
$M/\mathrm{H}]$ U(2.3,0.4)(-2.3,0.4) U(2.3,0.4)(-2.3,0.4)
$A_{V}$ U(0.0,4.0)(0.0,4.0) U(0.0,4.0)(0.0,4.0)
$z$ U(0.0,7.5)(0.0,7.5) U(0.0,7.5)(0.0,7.5)
@@
log($U$) 2.0-2.0
@@
  • $\dagger$ Formed stellar mass (integral of SFH). We also inferred the surviving stellar mass.
  • $\dagger\dagger$ Cosmic times at which 25%, 50%, and 75% of total stellar mass was formed, normalised to the age of the Universe.
@@
The output of the forward model were the fluxes in the 19 filters included in Table 1, including ACS filters and NIRCam wide-band and medium-band filters. To simulate realistic observational conditions, we modelled the flux uncertainties as a function of flux in each filter. We used the JADES segmentation map to identify all pixels classified as galaxies and, for each filter, we binned the pixel fluxes into five flux bins. Within each bin, we collected the flux uncertainties of the real observations and fit their distribution with a Gaussian. These fitted Gaussians, capturing the empirical uncertainty-flux relation, were then used to add noise to the simulated fluxes: for each simulated SED, the noise was sampled from the appropriate Gaussian based on its flux in each filter. This approach captures a broad range of uncertainties consistent with the observational data. Following the procedure described in Hahn & Melchior (2022), the network inputs included the fluxes with noise sampled from the distributions, as well as the standard deviations of the Gaussian distributions, to condition the inference by producing noise-aware posteriors.
@@
We transformed both fluxes and standard deviations to AB magnitudes, avoiding numerical instabilities in the training process caused by the wide dynamic range of possible flux values. We also took into account the $1\sigma$ depth limit in each filter $m_{i}^{b}$. We used the values obtained in Hainline et al. (2024) for the JADES mosaics. Although we repeated the calculations of the $1\sigma$ depth limits in the region were the galaxies were selected, we found negligible differences. To amortise this process, we used a mean value for the depth limit that was not dependent on the specific regions where the galaxies were located. If the simulated magnitude in a filter, $m_{i}$, for a CSP is fainter than that depth limit ($m_{i}>m_{\sigma_{i}}^{b}$), we transformed it to $m_{i}^{\prime}=100$ and and the noise in the input was $m_{\sigma_{i}}^{b}$. While $m_{i}^{\prime}=100$ is not physically meaningful (since it should be infinite), it is a convenient choice that avoids convergence problems in the calculations and allows for a clear identification of dropouts (i.e. cases where the signal is not detected in a given filter). These dropouts typically occur at bluer wavelengths than the Lyman break for high-z galaxies, although non-detections can also occur in redder filters, especially in low-S/N photometry. Identifying these dropouts could also ensure a correct redshift estimate, although we have not addressed it in this work. We obtained posterior distributions from the photometry to constrain the stellar population parameters included in Table 2 for CSPs, using the neural density estimator, as explained below.